How do you solve the systems 4x–7y=-8 and –3x+4y=6 by elimination
Answer by FrankM(1040)

4x–7y=-8
–3x+4y=6
Multiple first equation by 3 and second by 4
12x-21y=-24
-12x+16y=24
-5y=0
y=0
x=-2